OpenStudy (samthemoose):

Adam sells lemonade for $2 per cup and comics for $2.25 per issue. She earns $360 selling lemonade and comics. Write a linear model that shows the number of cups she sold to the number of comics she sold. If Adam sold 85 issues of comics, how many cups of lemonade did she sell?

OpenStudy (smhpak):

2c+2.25i=360

OpenStudy (peachesncream8):

\[2C+ 2.25i = 360\] \[C+I = 85\]

OpenStudy (peachesncream8):

\[c= i-85 \] then plug in c into the equation \[2(i-85)+ 2.25i=360\]

OpenStudy (peachesncream8):

then just combine like terms and solve